SURFACE CHAIN ARRANGEMENT OF POLYMER FILM STUDIED BY SURFACE-ENHANCED RAMAN SCATTERING

\underline{QI XUE}, Department of Polymer Science and Engineering, Nanjing University, Nanjing, 210093, P. R. China.

The chain arrangement on the surface of a polymer film was investigated by surface-enhanced Raman scattering using chemically deposited silver as the SERS activating layer. SERS measurement for samples of current interest, such polymeric liquid crystal were successfully performed without disturbing surface morphology. The results show that our SERS sampling technique by silver-coating layer is effective.